OpenStudy (anonymous):

which equation has a solution of d =145

OpenStudy (anonymous):

Do you have answer choices?

OpenStudy (anonymous):

yes 147-d=302 b.462-d=327 c. 517=d=382 d. 359=d=214

OpenStudy (anonymous):

Answer is D, 359=d-214 correct im I right or wrong. I first subtractracted 359 from 214 and gave me the answer of 145 and since d=145 I assume thats the answer can you please verify it .

OpenStudy (anonymous):

ah I'm sorry D: My laptop battery just died. I just lost everything I typed, let me try again

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Is the solution 145 or would you happen to have meant 135? I worked through all of them but only found a solution 135 and thought that that might be it..

OpenStudy (anonymous):

145

OpenStudy (anonymous):

A) 147 - d = 302 First, we need to get d alone so subtract 147 from both sides and you will have: -d = 155 There is still a negative 1 by the d, so we need to divide both sides by negative 1 to get rid of it. That will give us: d = -155 So A isn't the solution. Just showing you how to eliminate your answer choices, you'll have to do this to each one until you find one that has the right solution. For C) 517 - d = 382 Subtract both sides by 517 to get: -d = -135 Divide both sides by -1 to get: d = 135 That's the closest I could get to the solution. Let me work out D for you to see if you're right

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Can you help me on the steps tp get the answer for this other question please. maria drove 2 1/4 hours on a business trip. if she drove at a constant rate of 64 miles per hour, how far did she drive?

OpenStudy (anonymous):

D) 359 = d - 214 Add 214 to both sides. 573 = d We haven't tried B yet.. B) 462 - d = 327 Subtract 462 from both sides: -d = -135 Divide both sides by -1: d = 135.. wasn't that the same answer as C?? :O ahhh lol.. hold on @esshotwired @phi anyone know?

OpenStudy (anonymous):

If she drives 64 miles per hour, and she drove 2 1/4 hours, multiply them: 2 1/4 simplifies to 9/1 so now we just have 64/1 X 9/1 which equals 576 / 1 So she drove 576 miles. Hope that helps. : )